Farhan's brother is three times as old as him. in 12 years' time, the sum of their ages will be equal to 10 times farhan's present ages. find their present ages.
Answer by lwsshak3(11628)

Farhan's brother is three times as old as him. in 12 years' time, the sum of their ages will be equal to 10 times farhan's present ages. find their present ages.
***
let x=Farhan's present age
3x=His brother's present age
..
12 yrs from present:
x+12=Farhan's age
3x+12=Brother's age
x+12+3x+12=10x
4x+24=10x
6x=24
x=4
3x=12
Farhan's present age=4
Brother's present age=12